OpenStudy (anonymous):

What is the value of the following sum: 1!+2!+3!+4!+5!?

OpenStudy (anonymous):

Details and assumptions The number n!, read as n factorial, is equal to the product of all positive integers less than or equal to n. For example, 7!=7×6×5×4×3×2×1.

OpenStudy (anonymous):

1+2*1+3*2+4*3*2+5*4*3*2

OpenStudy (anonymous):

sorry it is not the correct answer

OpenStudy (anonymous):

What is the problem you are facing in it??

OpenStudy (anonymous):

i dont understand your question

OpenStudy (anonymous):

153

OpenStudy (anonymous):

Don't tell the answer directly, @SerikMB let him or her solve first..

OpenStudy (anonymous):

factorial n!=n*(n-1)*(n-2)*(n-3)... till n-x=1 ok?

OpenStudy (anonymous):

hey thanks i had made a mistake. sorry

OpenStudy (anonymous):

@waterineyes, i have shown it before, probably he got confused by addition

OpenStudy (anonymous):

\[1! = 1, \quad 2! = 2 \times 1 = 1, \quad 3! = 3 \times 2 \times 1 = 6\] \[4! = 4 \times 3 \times 2 \times 1 = 24, \quad 5! = 5 \times 4 \times 3 \times 2 \times 1 = 120\] Now just add them..

OpenStudy (anonymous):

1+2+6+24+120=153

OpenStudy (anonymous):

\[2! = 2 \times 1 = 2\]